# Continuous-Time Quantum Markov Chains And Discretizations Of $p$-Adic Schrödinger Equations: Comparisons And Simulations

### v2

@@ 8.11. The multiplication of distributions
We recall that the existence of the product $F\cdot G$ is equivalent to the existence of $\mathcal{F}\left[F\right]\ast\mathcal{F}\left[G\right]$. In addition, $\mathcal{F}\left[F\cdot G\right]=\mathcal{F}\left[F\right]\ast\mathcal{F}\left[G\right]$ and $\mathcal{F}\left[F\ast G\right]=\mathcal{F}\left[F\right]\cdot\mathcal{F}\left[G\right]$, see, e.g., [19, Section 7.5].
@@ 9.1. Preliminary results
where $\delta\left(x\right)$ is the Dirac distribution, see, e.g., [19, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. Notice that $\left|\widehat{J}(\left|\xi\right|_{p})\right|\leq 1$, for any $\xi\in\mathbb{Q}_{p}$.
@@ 9.2. $p$-Adic heat equations and Markov processes in the unit ball
We now show that $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)$, $x\in\mathbb{Z}_{p},t>0$, is a solution of the differential equation in (9.3). Indeed, by Lemmas 9.3, 9.5,
@@
in $\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right)$. Now, by Lemmas 9.2 and 9.3,
@@
We now review a result from ([40, Theorem 3.1]). We denote by $\mathcal{C}(\mathbb{Z}_{p})$ the $\mathbb{R}$-vector space of continuous functions on the unit ball. We recall that $\mathcal{D}\left(\mathbb{Z}_{p}\right)$ is dense in $\mathcal{C}(\mathbb{Z}_{p})$. We fix a time horizon $T\in\left(0,\infty\right]$, and consider the following initial value problem:
@@
The condition $J\left(\left|x\right|_{p}\right)\leq 1$, $x\in\mathbb{Z}_{p}$, was used in ([40, Theorem 3.1]) to assure that operator $\boldsymbol{J}$ satisfies the maximum principle. Consequently, here we do not need this condition. By ([40, Theorem 3.1]), there exists a probability measure $p_{t}\left(x,\cdot\right)$, $t\in\left[0,T\right]$, with $T=T(u_{0})$, $x\in\mathbb{Z}_{p}$, on the Borel $\sigma$-algebra $\mathcal{B}$ of $\mathbb{Z}_{p}$, such that the Cauchy problem (9.4) has a unique solution of the form
@@ 11. Appendix D: $p$-adic Schrödinger equations on the unit ball and quantum Markov chains
for any $\xi_{0}\in\mathbb{Z}_{p}$, $\xi\in\mathbb{Q}_{p}$, $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ is a distribution with support in $\mathbb{Z}_{p}$, see, e.g., [19, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. The explicit calculation of $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ can be carried out as in the case of the heat kernel, and one obtains that
@@ 11.1. CTQMCs
We now use Theorem 3.1 to construct a family of CTQMCs. We use all the notation introduced in Subsections 3.3, 3.4,
@@
We use all the notation introduced in Subsection $3.$ Set
@@
With the notation introduced in Propositions 11.1, 11.2.

### v3

@@ 9.11. The multiplication of distributions
We recall that the existence of the product $F\cdot G$ is equivalent to the existence of $\mathcal{F}\left[F\right]\ast\mathcal{F}\left[G\right]$. In addition, $\mathcal{F}\left[F\cdot G\right]=\mathcal{F}\left[F\right]\ast\mathcal{F}\left[G\right]$ and $\mathcal{F}\left[F\ast G\right]=\mathcal{F}\left[F\right]\cdot\mathcal{F}\left[G\right]$, see, e.g., [16, Section 7.5].
@@ 10.1. Preliminary results
where $\delta\left(x\right)$ is the Dirac distribution, see, e.g., [16, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. Notice that $\left|\widehat{J}(\left|\xi\right|_{p})\right|\leq 1$, for any $\xi\in\mathbb{Q}_{p}$.
@@ 10.2. $p$-Adic heat equations and Markov processes in the unit ball
We now show that $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)$, $x\in\mathbb{Z}_{p},t>0$, is a solution of the differential equation in (10.3). Indeed, by Lemmas 10.3, 10.5,
@@
in $\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right)$. Now, by Lemmas 10.2 and 10.3,
@@
We now review a result from ([29, Theorem 3.1]). We denote by $\mathcal{C}(\mathbb{Z}_{p})$ the $\mathbb{R}$-vector space of continuous functions on the unit ball. We recall that $\mathcal{D}\left(\mathbb{Z}_{p}\right)$ is dense in $\mathcal{C}(\mathbb{Z}_{p})$. We fix a time horizon $T\in\left(0,\infty\right]$, and consider the following initial value problem:
@@
The condition $J\left(\left|x\right|_{p}\right)\leq 1$, $x\in\mathbb{Z}_{p}$, was used in ([29, Theorem 3.1]) to assure that operator $\boldsymbol{J}$ satisfies the maximum principle. Consequently, here we do not need this condition. By ([29, Theorem 3.1]), there exists a probability measure $p_{t}\left(x,\cdot\right)$, $t\in\left[0,T\right]$, with $T=T(u_{0})$, $x\in\mathbb{Z}_{p}$, on the Borel $\sigma$-algebra $\mathcal{B}$ of $\mathbb{Z}_{p}$, such that the Cauchy problem (10.4) has a unique solution of the form
@@ 12. Appendix D: $p$-adic Schrödinger equations on the unit ball and quantum Markov chains
for any $\xi_{0}\in\mathbb{Z}_{p}$, $\xi\in\mathbb{Q}_{p}$, $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ is a distribution with support in $\mathbb{Z}_{p}$, see, e.g., [16, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. The explicit calculation of $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ can be carried out as in the case of the heat kernel, and one obtains that
@@ 12.1. CTQMCs
We now use Theorem 4.1 to construct a family of CTQMCs. We use all the notation introduced in Subsections 4.3, 4.4,
@@
We use all the notation introduced in Subsection 4.3. Set
@@
With the notation introduced in Propositions 12.1, 12.2.